# Learning an Image Editing Model without Image Editing Pairs

### v1

@@ 2 Related Works
Few-step diffusion models. Standard diffusion (or flow-matching) models require many sampling steps to generate high-quality images. Many prior works reduce the number of denoising steps for faster sampling by predicting larger denoising steps, including consistency models (Kim et al., 2024; Geng et al., 2024; Song et al., 2023a; Yang et al., 2024c; Song & Dhariwal, 2024; Lu & Song, 2025; Heek et al., 2024), shortcut models (Frans et al., 2024), meanflow (Geng et al., 2025), and inductive moment matching (Zhou et al., 2025). Another line distills a pre-trained multi-step teacher into a few-step student by matching ODE trajectories (Song et al., 2023a; Salimans & Ho, 2022; Geng et al., 2023), using adversarial loss (Sauer et al., 2024b; Kang et al., 2024; Yin et al., 2024a; Sauer et al., 2024a; Xu et al., 2024), or applying score distillation (Luo et al., 2023; Yin et al., 2024b; a; Zhou et al., 2024). In our framework, we adopt DMD (Yin et al., 2024b) as a distribution matching objective. This ensures that our few-step editing model’s output remains in the real-image manifold defined by the pre-trained text-to-image teacher, while using VLM-feedback to ensure the model follows the editing instructions.
@@ 4.1 Edit Instruction Dataset
Each dataset sample consists of a real image $\mathbf{y}$} as reference and an associated edit instruction, $\mathbf{c}$}. Following prior works (Liu et al., 2025b; Ye et al., 2025), we focus on several categories of local editing operations, such as Add, Replace, Remove, Adjust shape, Action, Stylization, Text editing, Color, Material, and Background change, as well as more free-form editing tasks such as Customization or Personalization (Gal et al., 2023; Ruiz et al., 2023; Kumari et al., 2023). Candidate instructions for each type are generated using Qwen2.5-32B VLM (Qwen-Team, 2025). Given an image-instruction pair, we further query the VLM to assess its validity and to suggest the caption, ${\mathbf{c}}^{{\mathbf{x}}}$, for the edited image. For the customization task, we restrict reference images to those showing a prominent central object, either filtered from a real image corpus or generated via the pretrained model (Tan et al., 2025; Kumari et al., 2025), and prompt the VLM to generate a caption that places the object in a novel background or context. In total, for local and free-form editing instructions, our dataset consists of $\sim 3$M and $\sim 600$K reference images, respectively. The input prompt to the Qwen2.5-32B VLM for each setup is shown in Appendix D.
@@ 4.2 Training Objective
${\mathbf{x}}_{\theta}^{0}$ $=\hat{{\mathbf{x}}}_{\theta}^{t}-t{\mathbf{v}}_{\theta},$ $\text{where }{\mathbf{v}}_{\theta}\equiv G_{\theta}(\hat{{\mathbf{x}}}_{\theta}^{t},t,{\mathbf{c}},{\mathbf{y}}),\quad\hat{{\mathbf{x}}}_{\theta}^{t}=(1-t)\hat{{\mathbf{x}}}_{\theta}^{0}+t\epsilon;\epsilon\sim\mathcal{N}(\bm{0},\bm{I}),t\sim(0,1).$
With the second step, the model is now trained on noisy intermediate states at timesteps determined by $t$, while being more efficient than a full backward unroll. In our method, we focus on few-step generation—specifically four steps—and restrict $t\in[0.25,0.5,0.75]$ in the second step. Few-step generator provides a better estimate of the denoised image, ${\mathbf{x}}^{0}_{\theta}$, at intermediate steps, which in turn enables effective VLM-based feedback. VLMs tend to give unreliable judgments when inputs are noisy or blurry (see Appendix B). This also enables faster inference and lowers training costs.
Figure 1: Method. We fine-tune a pretrained text-to-image model into a few-step image-editing model using differentiable VLM-feedback regarding edit success. In addition, we use distribution matching loss (DMD (Yin et al., 2024a)) to ensure output images remain in the natural image manifold.
@@
where $\ell^{(j)}_{a_{j}}$ is the logit corresponding to the token $\mathbf{X}_{a_{j}}$, $\sigma$ is the sigmoid function, and $p(a_{j})$ is the probability of correct answer, while restricting normalization to only the Yes and No tokens, which we observe to be more effective during training (Zhang et al., 2024). Computing this loss is relatively fast, as it only requires a single forward call to the VLM per question, as opposed to autoregressive token prediction.
@@
where ${\mathbf{c}}^{{\mathbf{x}}}$ is the text caption describing the noisy edited image ${\mathbf{x}}^{t}_{\theta}$ and ${\mathbf{v}}_{\text{real}}$, ${\mathbf{v}}_{\text{gen}}$ represents the predicted velocity from the teacher and a trainable auxiliary model, $A_{\phi}$ respectively. The auxiliary model is trained along with $G_{\theta}$ to learn the current output distribution of $G_{\theta}$ using a flow-based denoising objective. This loss ensures that the edited images not only satisfy the instruction but also remain faithful to the text conditioned distribution of real images modeled by the pretrained teacher.
@@ 4.3 Training details
1 $G_{\theta}\leftarrow\text{copyWeights}(G_{\text{init}})$;  $A_{\phi}\leftarrow\text{copyWeights}(G_{\text{init}})$
2
3// Warmup with identity loss
4 for $\text{step}=1$ to $N_{\text{warmup}}$ do
5 $({\mathbf{y}},{\mathbf{c}}^{\mathbf{y}})\sim\mathcal{X}$, $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$, $t\sim(0,1]$
6 ${\mathbf{x}}\leftarrow{\mathbf{y}}$
7 ${\mathbf{x}}^{t}\leftarrow(1-t){\mathbf{x}}+t\epsilon$
8 ${\mathbf{v}}_{\theta}\leftarrow G_{\theta}({\mathbf{x}}^{t},t,{\mathbf{y}},{\mathbf{c}}^{\mathbf{y}})$
@@ 7 ${\mathbf{x}}^{t}\leftarrow(1-t){\mathbf{x}}+t\epsilon$
10 $\theta_{G}\leftarrow\theta_{G}-\eta_{G}\nabla_{\theta_{G}}\,\mathcal{L}_{\text{id}}$.
11 end for
12
13
14// Main training loop
15 while train do
16 $\{({\mathbf{y}},{\mathbf{c}},{\mathbf{c}}^{{\mathbf{x}}})\}\sim\mathcal{X},\epsilon\sim\mathcal{N}(0,I),t\in[0.25,0.5,0.75,1]$
17 ${\mathbf{v}}_{\theta}\leftarrow G_{\theta}(\epsilon,t=1,{\mathbf{y}},{\mathbf{c}})$
18 ${\mathbf{x}}^{0}_{\theta}\leftarrow\epsilon-{\mathbf{v}}_{\theta}\quad$
19 if $t<1$ then
20
21    ${\mathbf{v}}_{\theta}\leftarrow G_{\theta}({\mathbf{x}}^{t}_{\theta},t,{\mathbf{y}},{\mathbf{c}});\quad{\mathbf{x}}^{t}_{\theta}\leftarrow(1-t){\mathbf{x}}^{0}_{\theta}\ +t\epsilon\;;\;\;\;\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$
22    ${\mathbf{x}}^{0}_{\theta}\leftarrow{\mathbf{x}}^{t}_{\theta}-t{\mathbf{v}}_{\theta}$
23   end if
24
25 $\text{Compute }\mathcal{L}_{\text{VLM}}$ // Eqn. 4
26 $\text{Compute }\nabla_{\theta}D_{KL}$ // Eqn. 5
27
28 $\theta_{G}\leftarrow\theta_{G}-\eta_{G}\lambda_{\text{vlm}}\nabla_{\theta_{G}}\mathcal{L}_{\text{VLM}}-\eta_{G}\lambda_{\text{dmd}}\nabla_{\theta}D_{KL}$
29
30 for $\text{local step}=1$ to $N_{\text{aux}}$ do
31
32    $\epsilon\sim\mathcal{N}(\bm{0},\bm{I}),\{({\mathbf{y}},{\mathbf{c}},{\mathbf{c}}^{{\mathbf{x}}})\}\sim\mathcal{X}$
33    ${\mathbf{x}}_{\theta}^{0}\leftarrow G_{\theta}(\epsilon,{\mathbf{y}},{\mathbf{c}})$ // edited image with backward unroll
34
35    ${\mathbf{x}}^{t}_{\theta}\leftarrow(1-t){\mathbf{x}}_{\theta}^{0}+t\epsilon\quad\epsilon\sim\mathcal{N}(\bm{0},\bm{I})\;t\in(0,1)$
36    ${\mathbf{v}}_{\phi}\leftarrow A_{\phi}({\mathbf{x}}^{t}_{\theta},{\mathbf{c}}^{\mathbf{x}})$
@@ 5.2 Free-form Editing: Customization
Evaluation metric. Here as well, we use the VIEScore evaluation with a similar Semantic Consistency (SC) score to evaluate identity and text alignment and the Perceptual Quality (PQ) score to measure realism, and the geometric mean of the two for the Overall score. We also report CLIPScore (Radford et al., 2021) and DINO (Oquab et al., 2023) similarity-based metrics in the Appendix.
@@ 5.3 Ablation
Training objective. We ablate our training objective across four settings: (1) using only distribution matching loss, (2) using only the VLM-based editing loss, (3) removing the identity-preservation loss $\mathcal{D}_{QA}$, and (4) replacing the binary-cross entropy loss (Eqn. 4) with standard cross-entropy over the full vocabulary. Results are shown below in Table 3. Training without the VLM-based loss and relying solely on distribution matching significantly degrades the model’s capabilities at following editing instructions. We observe that VLM-based loss is essential for maintaining consistency between input and edited images and for certain editing tasks like Removal (Figure 4 and Appendix Figure 5). However, only training with VLM-based loss leads to unrealistic outputs (Appendix Figure 6), and the training eventually diverges, as evidenced by the low overall score in Table 3, underscoring the need for DMD loss. In addition, using binary cross-entropy loss and having a question to check consistency between input and edited images improves the overall performance.
@@
The Appendix provides additional results and more detailed discussion of the method’s limitations.
@@ 6 Discussion and Limitations
Despite these promising results, our method has limitations. Without pixel-level supervision, edits may deviate from the input image in fine-grained details or fail to fully preserve subject identity. We show in Appendix C that adding a perceptual similarity loss (e.g., LPIPS (Zhang et al., 2018)) between input and edited images alleviates this to some extent, though often at the cost of editing quality. Another constraint for our method is the need to keep the VLM in GPU memory, introducing VRAM overhead. We expect ongoing advances in stronger and more efficient VLMs can help address this issue. Overall, our framework scales effectively with large unpaired datasets and highlights a path toward more flexible post-training of generative models for diverse downstream tasks.
Acknowledgment. We thank Gaurav Parmar, Maxwell Jones, and Ruihan Gao for their feedback and helpful discussions. This work was partly done while Nupur Kumari was interning at Adobe Research. The project was partly supported by Adobe Inc., the Packard Fellowship, the IITP grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project), NSF IIS-2239076, and NSF ISS-2403303.
@@ Appendix A Additional Comparison with Baseline Methods
Local image-editing. Here, we compare on another commonly adopted image-editing benchmark, ImgEdit (Ye et al., 2025) (Basic), which covers nine local editing types across diverse semantic categories with a total of $734$ samples. Quantitative results under their proposed GPT4o-based evaluation protocol are reported in Table 5, with VIEScore (Ku et al., 2024) results in Table 6. Consistent with the trend observed on GEdit-Bench, our method has better or on-par performance in the few-step setting and remains comparable to many of the multi-step baselines as well. Qualitative comparisons are shown in Figure 11, with additional examples on GEdit-Bench in Figure 10.

### v2

@@ 2 Related Works
Few-step diffusion models. Standard diffusion (or flow-matching) models require many sampling steps to generate high-quality images. Many prior works reduce the number of denoising steps for faster sampling by predicting larger denoising steps, including consistency models (Kim et al., 2024; Geng et al., 2024; Song et al., 2023a; Yang et al., 2024c; Song & Dhariwal, 2024; Lu & Song, 2025; Heek et al., 2024; Frans et al., 2024; Zhou et al., 2025), shortcut models (Frans et al., 2024), meanflow (Geng et al., 2025), and inductive moment matching (Zhou et al., 2025). Another line distills a pre-trained multi-step teacher into a few-step student by matching ODE trajectories (Song et al., 2023a; Salimans & Ho, 2022; Geng et al., 2023), using adversarial loss (Sauer et al., 2024b; Kang et al., 2024; Yin et al., 2024a; Sauer et al., 2024a; Xu et al., 2024), or applying score distillation (Luo et al., 2023; Yin et al., 2024b; a; Zhou et al., 2024). In our framework, we adopt DMD (Yin et al., 2024b) as a distribution matching objective. This ensures that our few-step editing model’s output remains in the real-image manifold defined by the pre-trained text-to-image teacher, while using VLM-feedback to ensure the model follows the editing instructions.
@@ 4.1 Edit Instruction Dataset
Each dataset sample consists of a real image $\mathbf{y}$} as reference and an associated edit instruction, $\mathbf{c}$}. Following prior works (Liu et al., 2025b; Ye et al., 2025), we focus on several categories of local editing operations, such as Add, Replace, Remove, Adjust shape, Action, Stylization, Text editing, Color, Material, and Background change, as well as more free-form editing tasks such as Customization or Personalization (Gal et al., 2023; Ruiz et al., 2023; Kumari et al., 2023). Candidate instructions for each type are generated using Qwen2.5-32B VLM (Qwen-Team, 2025). Given an image-instruction pair, we further query the VLM to assess its validity and to suggest the caption, ${\mathbf{c}}^{{\mathbf{x}}}$, for the edited image. For the customization task, we restrict reference images to those showing a prominent central object, either filtered from a real image corpus or generated via the pretrained model (Tan et al., 2025; Kumari et al., 2025), and prompt the VLM to generate a caption that places the object in a novel background or context. In total, for local and free-form editing instructions, our dataset consists of $\sim 3$M and $\sim 600$K reference images, respectively. The input prompt to the Qwen2.5-32B VLM for each setup is shown in the Appendix D.
@@ 4.2 Training Objective
${\mathbf{x}}_{\theta}^{0}$ $=\hat{{\mathbf{x}}}_{\theta}^{t}-t{\mathbf{v}}_{\theta},$ $\text{where }{\mathbf{v}}_{\theta}\equiv G_{\theta}(\hat{{\mathbf{x}}}_{\theta}^{t},t,{\mathbf{c}},{\mathbf{y}}),\quad\hat{{\mathbf{x}}}_{\theta}^{t}=(1-t)\hat{{\mathbf{x}}}_{\theta}^{0}+t\epsilon;\epsilon\sim\mathcal{N}(\bm{0},\bm{I}),t\sim(0,1),$
With the second step, the model is now trained on noisy intermediate states at timesteps determined by $t$, while being more efficient than a full backward unroll. In our method, we focus on few-step generation—specifically four steps—and restrict $t\in[t_{1},t_{2},t_{3}]$, a fixed schedule, in the second step. Few-step generator provides a better estimate of the denoised image, ${\mathbf{x}}^{0}_{\theta}$, at intermediate steps, which in turn enables effective VLM-based feedback. VLMs tend to give unreliable judgments when inputs are noisy or blurry (see Appendix B). This also enables faster inference and lowers training costs.
Figure 1: Method. We fine-tune a pretrained text-to-image model into a few-step image-editing model using differentiable VLM-feedback regarding edit success and distribution matching loss (DMD (Yin et al., 2024a)) for ensuring output images remain in the natural image manifold. Given the edit instruction and condition image, we predict the edited image starting from noise. During training, we randomly sample the few-shot timestep (Line 3 of Algorithm 1), and perform a two-step diffusion unrolling to predict the edited image for the intermediate timestep as shown here. The loss is backpropagated through the two diffusion sampling steps.
@@
where $\ell^{(j)}_{a_{j}}$ is the logit corresponding to the token $\mathbf{X}_{a_{j}}$, $\sigma$ is the sigmoid function, and $p(a_{j})$ is the probability of correct answer, while restricting normalization to only the Yes and No tokens, which we observe to be more effective during training (Zhang et al., 2024). Computing this loss is relatively fast, as it requires a single forward call to the VLM per question, as opposed to autoregressive prediction.
@@
where ${\mathbf{c}}^{{\mathbf{x}}}$ is the text caption describing the noisy edited image ${\mathbf{x}}^{t}_{\theta}$ and ${\mathbf{v}}_{\text{real}}$, ${\mathbf{v}}_{\text{gen}}$ represents the predicted velocity from the teacher and a trainable auxiliary model, $A_{\phi}$ respectively. The auxiliary model is trained along with $G_{\theta}$ to learn the current output distribution of $G_{\theta}$ using a flow-based denoising objective. This loss ensures that the edited images not only satisfy the instruction but also remain faithful to the text-conditioned distribution of real images modeled by the pretrained teacher.
@@ 4.3 Training details
$G_{\theta}\leftarrow\text{copyWeights}(G_{\text{init}})$;  $A_{\phi}\leftarrow\text{copyWeights}(G_{\text{init}})$
// Warmup with identity loss
for $\text{step}=1$ to $N_{\text{warmup}}$ do
 $({\mathbf{y}},{\mathbf{c}}^{\mathbf{y}})\sim\mathcal{X}$, $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$, $t\sim(0,1]$
 ${\mathbf{x}}\leftarrow{\mathbf{y}}$
 ${\mathbf{x}}^{t}\leftarrow(1-t){\mathbf{x}}+t\epsilon$
 ${\mathbf{v}}_{\theta}\leftarrow G_{\theta}({\mathbf{x}}^{t},t,{\mathbf{y}},{\mathbf{c}}^{\mathbf{y}})$
@@
 $\theta_{G}\leftarrow\theta_{G}-\eta_{G}\nabla_{\theta_{G}}\,\mathcal{L}_{\text{id}}$.
end for
// Main training loop
while train do
 $\{({\mathbf{y}},{\mathbf{c}},{\mathbf{c}}^{{\mathbf{x}}})\}\sim\mathcal{X},\epsilon\sim\mathcal{N}(0,I),t\in[t_{1},t_{2},t_{3},t_{4}=1]$
 ${\mathbf{v}}_{\theta}\leftarrow G_{\theta}(\epsilon,t=1,{\mathbf{y}},{\mathbf{c}})$
 ${\mathbf{x}}^{0}_{\theta}\leftarrow\epsilon-{\mathbf{v}}_{\theta}\quad$
 if $t<1$ then
    ${\mathbf{v}}_{\theta}\leftarrow G_{\theta}({\mathbf{x}}^{t}_{\theta},t,{\mathbf{y}},{\mathbf{c}});\quad{\mathbf{x}}^{t}_{\theta}\leftarrow(1-t){\mathbf{x}}^{0}_{\theta}\ +t\epsilon\;;\;\;\;\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$
    ${\mathbf{x}}^{0}_{\theta}\leftarrow{\mathbf{x}}^{t}_{\theta}-t{\mathbf{v}}_{\theta}$
   end if
 $\text{Compute }\mathcal{L}_{\text{VLM}}$ // Eqn. 4
 $\text{Compute }\nabla_{\theta}D_{KL}$ // Eqn. 5
 $\theta_{G}\leftarrow\theta_{G}-\eta_{G}\lambda_{\text{vlm}}\nabla_{\theta_{G}}\mathcal{L}_{\text{VLM}}-\eta_{G}\lambda_{\text{dmd}}\nabla_{\theta}D_{KL}$
 for $\text{local step}=1$ to $N_{\text{aux}}$ do
    $\epsilon\sim\mathcal{N}(\bm{0},\bm{I}),\{({\mathbf{y}},{\mathbf{c}},{\mathbf{c}}^{{\mathbf{x}}})\}\sim\mathcal{X}$
    ${\mathbf{x}}_{\theta}^{0}\leftarrow G_{\theta}(\epsilon,{\mathbf{y}},{\mathbf{c}})$ // edited image with backward unroll
    ${\mathbf{x}}^{t}_{\theta}\leftarrow(1-t){\mathbf{x}}_{\theta}^{0}+t\epsilon\quad\epsilon\sim\mathcal{N}(\bm{0},\bm{I})\;t\in(0,1)$
    ${\mathbf{v}}_{\phi}\leftarrow A_{\phi}({\mathbf{x}}^{t}_{\theta},{\mathbf{c}}^{\mathbf{x}})$
@@ 5.2 Free-form Editing: Customization
Evaluation metric. Here as well, we use the VIEScore evaluation with a similar Semantic Consistency (SC) score to evaluate identity and text alignment, a Perceptual Quality (PQ) score to measure realism, and the geometric mean of the two for the Overall score. We also report CLIPScore (Radford et al., 2021) and DINO (Oquab et al., 2023) similarity-based metrics in the Appendix.
@@ 5.3 Ablation
Training objective. We ablate our training objective across four settings: (1) using only distribution matching loss, (2) using only the VLM-based editing loss, (3) removing the identity-preservation question from $\mathcal{D}_{QA}$, and (4) replacing the binary-cross entropy loss (Eqn. 4) with standard cross-entropy over the full vocabulary. Results are shown below in Table 3. Training without the VLM-based loss and relying solely on distribution matching significantly degrades the model’s capabilities at following editing instructions. We observe that VLM-based loss is essential for maintaining consistency between input and edited images and for certain editing tasks like Removal (Figure 4 and Appendix Figure 5). However, only training with VLM-based loss leads to unrealistic outputs (Appendix Figure 6), and the training eventually diverges, as evidenced by the low overall score in Table 3, underscoring the need for DMD loss. In addition, using binary cross-entropy loss and having a question to check consistency between input and edited images improves the overall performance.
@@
The Appendix provides additional results and a more detailed discussion of the method’s limitations.
@@ 6 Discussion and Limitations
Despite these promising results, our method has limitations. Without pixel-level supervision, edits may deviate from the input image in fine-grained details or fail to fully preserve subject identity. We show in Appendix C that adding a perceptual similarity loss (e.g., LPIPS (Zhang et al., 2018)) between input and edited images alleviates this to some extent, though often at the cost of editing quality. Another concern for our method is that it is directly tied to the VLM’s capabilities and its biases. Additionally, the requirement to keep it in GPU memory introduces VRAM overhead. We expect ongoing advances in stronger and more efficient VLMs can help address this issue. Overall, our framework scales effectively with large unpaired datasets and highlights a path toward more flexible post-training of generative models for diverse downstream tasks.
Acknowledgment. We thank Gaurav Parmar, Maxwell Jones, and Ruihan Gao for their feedback and helpful discussions. This work was partly done while Nupur Kumari was interning at Adobe Research. Sheng-Yu Wang is supported by the Google PhD Fellowship. The project was partly supported by Adobe Inc., the Packard Fellowship, the IITP grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project), NSF IIS-2239076, and NSF ISS-2403303.
@@ Appendix A Additional Comparison with Baseline Methods
Human preference study. We perform a pairwise human comparison study on GEdit-Bench using Amazon Mechanical Turk. To measure editing success (similar to SC score in VIEScore), following Imagic (Kawar et al., 2023), we show the reference image in the middle and the images edited by our method and a baseline method, on the left and right randomly, and ask the user: “Which of the two images (Left or Right) matches with the editing instruction better, while preserving the details from the input reference image shown in the middle?”. We collect $500$ responses per pairwise comparison and show the results in Table 5. Under the few-step setting, our method outperforms Turbo-Edit, marginally preferred over FLUX.1-Kontext, Step-1X Edit, and has a random chance when compared with Qwen-Image-Edit. This is consistent with the VIEScore evaluation in Table 1. Similar to editing success, to measure perceptual quality (similar to PQ score in VIEScore), we show the images edited by our method and a baseline method on the left and right randomly, and ask the user: “Which image do you think has better overall quality and photorealism?”. As Table 5 shows, here our method is preferred over all the baselines, also consistent with the VIEScore evaluation.
Local image-editing. Here, we compare on another commonly adopted image-editing benchmark, ImgEdit (Ye et al., 2025) (Basic), which covers nine local editing types across diverse semantic categories with a total of $734$ samples. Quantitative results under their proposed GPT4o-based evaluation protocol are reported in Table 6, with VIEScore (Ku et al., 2024) results in Table 7. Consistent with the trend observed on GEdit-Bench, our method has better or on-par performance in the few-step setting and remains comparable to many of the multi-step baselines as well. We also include a comparison on TEDBench (Kawar et al., 2023), which focuses primarily on non-rigid and background edit-types. Table 8 shows the quantitative comparison with similar trends as GEdit-Bench and ImgEdit (Basic). Qualitative comparisons are shown in Figure 14 and 16 for ImgEdit (Basic) and TEDBench, respectively. We also show additional qualitative samples on GEdit-Bench in Figure 13.
@@
fal
Table 5: Human evaluation on GEdit-Bench under few-step sampling setting. We report the percentage of users who prefer our method over each baseline. For editing success, we outperform Turbo-Edit and remain competitive with other baselines, with a slight preference over FLUX.1-Kontext and Step-1X Edit and random chance with Qwen-Image-Edit. This trend is consistent with the VIEScore evaluation in Table 1. The standard error for all is within $\pm 4\%$.
Ours vs Turbo-Edit FLUX.1-Kontext Step-1X Edit Qwen-Image-Edit
SC preference 63.19 $\%$ 53.35 $\%$ 56.46 $\%$ 48.44 $\%$
PQ preference 73.92 $\%$ 79.91 $\%$ 80.39 $\%$ 71.98 $\%$
@@
Table 8: VIEScore evaluation on TEDBench. Our method performs on par or better than baselines under the few-step setting and remains competitive to baselines with multi-step sampling. All numbers reported in $\times 10$
[…]
FLUX.1-Kontext 12B 28 6.54 6.77 5.97
Step-1X Edit v1.1 12B 28 6.86 7.62 6.67
QwenImage Edit 20B 50 7.90 7.88 7.58
QwenImage Edit 20B 4 6.20 5.32 5.44
FLUX.1-Kontext 12B 4 5.84 4.96 4.92
Step-1X Edit v1.1 12B 4 6.24 5.90 5.39
NP-Edit (Ours) 2B 4 6.96 8.28 6.98